# Fast measurements and multiqubit gates in dual species atomic arrays

### v1

@@ I Introduction
The level scheme for the state mapping between the Cs ancilla and Rb measuring atoms is shown in Fig. 2. It is based on a modified form of the Rydberg controlled state transfer described in [17], with the significant difference that the transfer is performed by a microwave field instead of an optical laser field. The protocol operates as follows. Rb atoms are prepared in state |$\ket{0}_{\rm m}$ which is dark to the interrogating laser. The Cs ancilla is either in state |$\ket{0}_{\rm a}$ and remains there, or in state |$\ket{1}_{\rm a}$ and is transferred to the Rydberg state |$\ket{r}_{\rm a}$. A microwave $\pi$ pulse resonant with the transition |$\ket{0}_{\rm m}\to\ket{1}_{\rm m}$ is applied to the Rb atoms. State |$\ket{1}_{\rm m}$ is also coupled to a Rydberg state |$\ket{r}_{\rm m}$ by a non-resonant laser leading to a dynamic Stark shift of |$\ket{1}_{\rm m}$ that depends on whether Cs atom is in non-interacting state |$\ket{0}_{\rm a}$ or in the strongly-interacting Rydberg state |$\ket{r}_{\rm a}$. In the former case, the Stark shift is large making the microwave pulse non-resonant and thus suppressing the transfer |$\ket{0}_{\rm m}\to\ket{1}_{\rm m}$. In the latter case, the Rydberg level |$\ket{r}_{\rm m}$ is largely shifted by the strong interaction with the Cs atom, which in turn reduces the Stark shift of |$\ket{1}_{\rm m}$ making the transition |$\ket{0}_{\rm m}\to\ket{1}_{\rm m}$ resonant with the microwave pulse, transferring the Rb atoms to |$\ket{1}_{\rm m}$. This implements the mapping described in Eqs. (1).
After the state mapping is performed, the Rb measuring atoms are subjected to an interrogating laser acting on the cycling transition |$\ket{1}_{\mathrm{m}}\to\ket{f}_{\mathrm{m}}$ and their state is recorded. The measuring atoms optically pumped to |$\ket{1}_{\rm m}$ can then be rapidly returned to the initial state |$\ket{0}_{\rm m}$ by a microwave $\pi$ pulse. The Cs ancillae also need to be reset. This might appear to be problematic since scattering of repump photons from the Cs ancillae could cause dephasing of nearby Cs data qubits; however, the measurement results tell us which state the Cs ancillae are in and local $\pi$ pulses on the qubit transition can be used to reset the ancillae as needed, so optical pumping of the Cs atoms is not required.
@@ II Dynamics of the measuring atoms
In Fig. 4 we show the dynamics of population transfer between the states |$\ket{0}_{\rm m}$ and |$\ket{1}_{\rm m}$ of the measuring Rb atoms, as obtained from the numerical solution of the Schrödinger equation with the full Hamiltonian (2). The numerical results show that when the Cs atom is in state |$\ket{0}_{\rm a}$ the transfer of the measurement atoms is suppressed to less than $10^{-4}$. When the Cs atom is in state |$\ket{1}_{\rm a}$ the measurement atom transfer is resonant and succeeds with probability greater than 80%. The combination of the very high fidelity of suppression (ancilla in |$\ket{0}_{\rm a}$) and high probability of multi-atom transfer (ancilla in |$\ket{1}_{\rm a}$) leads to low measurement error with a short integration time, as is detailed in Sec. III below.
@@
Another potential concern is the possible ionization of Rydberg states by the 6.8 GHz microwave field used to drive the |$\ket{0}_{\rm m}\rightarrow\ket{1}_{\rm m}$ transition. Ionization of non-hydrogenic Rydberg states can occur at a relatively low threshold field of $E_{\rm at}/(3n^{5})$ with $E_{\rm at}=5.1\times 10^{11}\leavevmode\nobreak\ \rm V/m$ [20]. However, for a circularly polarized microwave field, as is used to drive the |$\ket{0}_{\rm m}\rightarrow\ket{1}_{\rm m}$ transition, the ionization threshold is $E_{\rm ionize}=E_{\rm at}/(16n^{4})$ [21]. At a microwave Rabi frequency of $\Omega_{\mathrm{max}}=2\pi\times 0.2\;$MHz the field strength is $E_{\rm max}=3500\>$V/m which is much less than the $n=48$ ionization field of $E_{\rm ionize}=6050\>$V/m. This field strength corresponds to about 1.6 W of microwave power in an area of $1\leavevmode\nobreak\ \rm cm^{2}$ and the power could be increased by more than a factor of two before reaching the ionization thresold. Alternatively, for faster transfer between the ground-state hyperfine sublevels, we can use two-photon optical Raman transition [22, 23].
@@ III Measurement and error correction cycle time
When the atom is in the dark state |$\ket{0}_{\mathrm{m}}$, it does not scatter any photons but photoelectron counts accumulate due to detector dark rate and background light scattering. Assuming negligible background light scattering, we can write for the mean background count $b(t_{\rm m})=b_{0}t_{\rm m}$ with $b_{0}$ due to detector dark counts. When the atoms are in the bright state |$\ket{1}_{\mathrm{m}}$, the mean signal is $s(t_{\rm m})=kq(t_{\rm m})+b(t_{\rm m})$ for $k$ measuring atoms. After the exposure, photoelectrons are converted to camera counts by digitizing the voltage of the well such that the ratio of camera counts to photoelectrons is given by a calibration factor $\kappa$.
@@
where $P_{\rm dark},P_{\rm bright}$ are the probabilities that the atom is in the dark or bright states. For given values of $t_{\rm m}$, $q(t_{\rm m})$, $b(t_{\rm m})$, $\sigma$, and $\kappa$, there is an optimum choice of $n_{\rm t}$ that minimizes the error. Figure 5 shows measurement error as a function of integration time for up to $k=5$ atoms. The calculated curves assume the background noise is entirely due to camera electronics and not background optical scattering. This is the case using a geometry with readout beams propagating in a plane perpendicular to the axis of the collection lens [3]. We see that with 5 measurement atoms an infidelity less than $10^{-4}$ is possible with an integration time $t_{\mathrm{m}}<5\leavevmode\nobreak\ \mu\rm s$.
@@ IV Multiqubit ${\sf CNOT}_{k}$ gate
To satisfy these conditions, we assume the Rydberg dressing laser with $\Omega_{r}/2\pi=10\>$MHz and $\Delta_{r}/2\pi=25\>$MHz with a Cs-Rb interaction strength $B_{0j}\gtrsim 4\Delta_{r}$, leading to $0.2<\mu\beta/2\pi<1\>$MHz. With $\mu=3$ and $\beta/2\pi=0.18\>$MHz, we set the pulse duration $\tau=7\>\mu$s and its peak Rabi frequency $\Omega_{\mathrm{max}}/2\pi=0.6\>$MHz so that the exponent in the Landau-Zener formula is $\simeq 5.8$ and the nonadiabatic transition probability is less than $3\times 10^{-3}$. Such a large Rabi frequency for the magnetic dipole transition |$\ket{0}_{\mathrm{m}}\leftrightarrow\ket{1}_{\mathrm{m}}$ of the Rb atoms would require the microwave field amplitude $E_{\rm max}\gtrsim 10^{4}\>$/m which is much stronger than the ionization field for the Rydberg states |$\ket{r}_{\mathrm{a}}\equiv\ket{64s_{1/2},m_{j}=1/2}$ of Cs and |$\ket{r}_{\mathrm{m}}\equiv\ket{60s_{1/2},m_{j}=1/2}$ of Rb that we now consider. Instead, we assume a two-photon optical Raman transition between the ground-state hyperfine sublevels |$\ket{0}_{\rm m}\equiv\ket{5s_{1/2},f=1,m_{f}=0}$ and |$\ket{1}_{\rm m}\equiv\ket{5s_{1/2},f=2,m_{f}=0}$ via the intermediate non-resonant state |$\ket{5p_{1/2},f=1,m_{f}=1}$.
@@
where $\phi_{0}=2.03\>$rad is the single-atom dynamical phase accumulated during the transfer. We simulate the dynamics of the complete system, composed of one control qubit and $k=1,2,3,4$ target qubits, governed by the Hamiltonian (2). The resulting fidelity $\mathcal{F}=|\braket{\mathrm{GHZ}}{\Psi(\tau)}|^{2}$ of the GHZ (or Bell, for $k=1$) state is shown in the inset of Fig. 7. We obtain $\mathcal{F}\geq 0.98$ for all $k\leq 4$. The main factor that reduces the fidelity is the decay of the Rydberg state |$\ket{r}_{\mathrm{a}}$ of the Cs atom during the long gate time $\tau$. Then the fidelity decreases with increasing the number of target atoms due to the decay of the Rydberg state |$\ket{r}_{\mathrm{m}}$ with which we dress the qubit state |$\ket{1}_{\mathrm{m}}$ and therefore lose during the transfer some population from that state, $\Delta\braket{\sigma_{11}^{(j)}}\lesssim\frac{1}{2}\gamma_{r}\tau\frac{|\Omega_{r}|^{2}}{4\Delta_{r}^{2}}$. Again, going to higher Rydberg states for both Cs and Rb, and/or using the longer-lived $nP$ states, will increase the gate fidelity, and multiqubit Rydberg gates with error probability of $<1\%$ appear feasible under realistic experimental conditions.
@@ Appendix A Rydberg state parameters and calculation methodology
Atomic pair state interaction strengths were calculated with ARC [51, 50]. Pair states were calculated in a $\theta=\frac{\pi}{2},\phi=0$ geometry, with $m_{j}=1/2$. The interaction Hamiltonian was constructed from a range of states specified by $\Delta n=\pm 2,\Delta l=\pm 2$. For each Hamiltonian (Rb-Rb, Cs-Cs, and Rb-Cs), 250 eigenstates were extracted for the inter-atomic separation distances $r=1.5-6\;\mu$m. For each case, the arithmetic mean of the pair interaction curves, weighted by the overlap of the eigenstate with the target state $g=|\braket{\psi_{\rm pair-state}}{\psi_{\rm target}}|^{2}$, was fitted by a function of the form $V(r)=\frac{C_{3}}{r^{3}}+\frac{C_{6}}{r^{6}}$. Only states with $g\geq 0.4$ for Rb-Rb and Cs-Cs and $g\geq 0.01$ for Rb-Cs were included. The fits were performed over the interval $3\leavevmode\nobreak\ \mu$m $\leq r\leq 6\leavevmode\nobreak\ \mu$m, and results are presented in Fig. 8 and Table 2.

### v2

@@ I Introduction
The level scheme for the state mapping between the Cs ancilla and Rb measuring atoms is shown in Fig. 2. The mapping is based on a modified form of Rydberg controlled state transfer [17, 18]. The protocol operates as follows. Rb atoms are prepared in state |$\ket{0}_{\rm m}$ which is dark to the interrogating laser. The Cs ancilla is either in state |$\ket{0}_{\rm a}$ and remains there, or in state |$\ket{1}_{\rm a}$ and is transferred to the Rydberg state |$\ket{r}_{\rm a}$. A microwave $\pi$-pulse resonant with the transition |$\ket{0}_{\rm m}\to\ket{1}_{\rm m}$ is applied to the Rb atoms. State |$\ket{1}_{\rm m}$ is also coupled to a Rydberg state |$\ket{r}_{\rm m}$ by a non-resonant laser leading to a dynamic Stark shift of |$\ket{1}_{\rm m}$ that depends on whether Cs atom is in non-interacting state |$\ket{0}_{\rm a}$ or in the strongly-interacting Rydberg state |$\ket{r}_{\rm a}$. In the former case, the Stark shift is large making the microwave pulse non-resonant and thus suppressing the transfer |$\ket{0}_{\rm m}\to\ket{1}_{\rm m}$. In the latter case, the Rydberg level |$\ket{r}_{\rm m}$ is largely shifted by the strong interaction with the Cs atom, which in turn reduces the Stark shift of |$\ket{1}_{\rm m}$ making the transition |$\ket{0}_{\rm m}\to\ket{1}_{\rm m}$ resonant with the microwave pulse, transferring the Rb atoms to |$\ket{1}_{\rm m}$. This implements the mapping described in Eqs. (1).
After the state mapping is performed, the Rb measuring atoms are subjected to an interrogating laser acting on the cycling transition |$\ket{1}_{\mathrm{m}}\to\ket{f}_{\mathrm{m}}$ and their state is recorded. The measuring atoms optically pumped to |$\ket{1}_{\rm m}$ can then be rapidly returned to the initial state |$\ket{0}_{\rm m}$ by another microwave $\pi$-pulse. The Cs ancillae also need to be reset. This might appear to be problematic since scattering of repump photons from the Cs ancillae could cause dephasing of nearby Cs data qubits; however, the measurement results tell us which state the Cs ancillae are in and local $\pi$ pulses on the qubit transition can be used to reset the ancillae as needed, so optical pumping of the Cs atoms is not required.
@@ II Dynamics of the measuring atoms
In Fig. 4 we show the dynamics of population transfer between the states |$\ket{0}_{\rm m}$ and |$\ket{1}_{\rm m}$ of the measuring Rb atoms, as obtained from the numerical solution of the Schrödinger equation $i\partial_{t}\ket{\Psi}=H\ket{\Psi}$ for the wavevector |$\ket{\Psi}$ of the system governed by the non-Hermitian Hamiltonian (2), assuming that decay of the Rydberg states of Cs and Rb atoms leads to the loss of population to states outside the computational basis. The numerical results show that when the Cs atom is in state |$\ket{0}_{\rm a}$ the transfer of the measurement atoms is suppressed to less than $10^{-4}$. When the Cs atom is in state |$\ket{1}_{\rm a}$ the measurement atom transfer is resonant and succeeds with probability greater than 80%. The combination of the very high fidelity of suppression (ancilla in |$\ket{0}_{\rm a}$) and high probability of multi-atom transfer (ancilla in |$\ket{1}_{\rm a}$) leads to low measurement error with a short integration time, as detailed in Sec. III below.
@@
Another potential concern is the possible ionization of Rydberg states by the 6.8 GHz microwave field used to drive the |$\ket{0}_{\rm m}\rightarrow\ket{1}_{\rm m}$ transition. Ionization of non-hydrogenic Rydberg states can occur at a relatively low threshold field of $E_{\rm at}/(3n^{5})$ with $E_{\rm at}=5.1\times 10^{11}\leavevmode\nobreak\ \rm V/m$ [21]. However, for a circularly polarized microwave field, as is used to drive the |$\ket{0}_{\rm m}\rightarrow\ket{1}_{\rm m}$ transition, the ionization threshold is $E_{\rm ionize}=E_{\rm at}/(16n^{4})$ [22]. At a microwave Rabi frequency of $\Omega_{\mathrm{max}}=2\pi\times 0.2\;$MHz the field strength is $E_{\rm max}=3500\>$V/m which is much less than the $n=48$ ionization field of $E_{\rm ionize}=6050\>$V/m. This field strength corresponds to about 1.6 W of microwave power in an area of $1\leavevmode\nobreak\ \rm cm^{2}$ and the power could be increased by more than a factor of two before reaching the ionization thresold. Alternatively, for faster transfer between the ground-state hyperfine sublevels, we can use two-photon optical Raman transition [23, 18, 24].
@@ III Measurement and error correction cycle time
When the atom is in the dark state |$\ket{0}_{\mathrm{m}}$, it does not scatter any photons but photoelectron counts accumulate due to detector dark rate and background light scattering. Assuming negligible background light scattering, we can write for the mean background count $b(t_{\rm m})=b_{0}t_{\rm m}$ with $b_{0}$ due to detector dark counts. When the atoms are in the bright state |$\ket{1}_{\mathrm{m}}$, the mean detected signal is $s(t_{\rm m})=kq(t_{\rm m})+b(t_{\rm m})$ for $k$ measuring atoms. After the exposure, photoelectrons are converted to camera counts by digitizing the voltage of the well such that the ratio of camera counts to photoelectrons is given by a calibration factor $\kappa$. The critical factor for reaching high measurement fidelity is the ratio of signal to background counts, $s(t_{\rm m})/b(t_{\rm m})=k[q(t_{\rm m})/b(t_{\rm m})]+1$, which, in the limit of small background $b(t_{\rm m})\ll q(t_{\rm m})$, is proportional to the number of measurement atoms $k$. We can equivalently state that the measurement time $t_{\rm m}$ to reach a desired fidelity scales inversely with $k$.
@@
where $P_{\rm dark},P_{\rm bright}$ are the probabilities that the atom is in the dark or bright states. For given values of $t_{\rm m}$, $q(t_{\rm m})$, $b(t_{\rm m})$, $\sigma$, and $\kappa$, there is an optimum choice of $n_{\rm t}$ that minimizes the error. Figure 5 shows measurement error as a function of integration time for up to $k=5$ atoms. The calculated curves assume the background noise is entirely due to camera electronics and not background optical scattering. This is the case using a geometry with readout beams propagating in a plane perpendicular to the axis of the collection lens [3]. We see that with 5 measurement atoms an infidelity less than $10^{-4}$ is possible with an integration time $t_{\mathrm{m}}<5\leavevmode\nobreak\ \mu\rm s$. It should be noted that the required integration time for a desired measurement fidelity depends on the photoelectron generation rate which in turn is a function of the intensity and detuning of the readout light, as given by Eq. (6). The readout light parameters also determine the equilibrium atomic temperature and for the parameters of Fig. 5 a simplified one-dimensional theory [26] predicts an equilibrium atom temperature of $510\leavevmode\nobreak\ \mu\rm K$. However, for measurement times of $t_{\rm m}\leq 25\leavevmode\nobreak\ \mu\rm s$ the scattering induced heating amounts to $\leq 40\leavevmode\nobreak\ \mu\rm K$ in each of three dimensions. Thus fast measurements on a $\mu\rm s$ timescale are consistent with low atom loss using trap depths of a few hundred $\mu\rm K$, as has been demonstrated recently [27].
@@ IV Multiqubit ${\sf CNOT}_{k}$ gate
To satisfy these conditions, we assume the Rydberg dressing laser with $\Omega_{r}/2\pi=10\>$MHz and $\Delta_{r}/2\pi=25\>$MHz with a Cs-Rb interaction strength $B_{0j}\gtrsim 4\Delta_{r}$, leading to $0.2<\mu\beta/2\pi<1\>$MHz. With $\mu=3$ and $\beta/2\pi=0.18\>$MHz, we set the pulse duration $\tau=7\>\mu$s and its peak Rabi frequency $\Omega_{\mathrm{max}}/2\pi=0.6\>$MHz so that the exponent in the Landau-Zener formula is $\simeq 5.8$ and the nonadiabatic transition probability is less than $3\times 10^{-3}$. Such a large Rabi frequency for the magnetic dipole transition |$\ket{0}_{\mathrm{m}}\leftrightarrow\ket{1}_{\mathrm{m}}$ of the Rb atoms would require the microwave field amplitude $E_{\rm max}\gtrsim 10^{4}\>$/m which is much stronger than the ionization field for the Rydberg states |$\ket{r}_{\mathrm{a}}\equiv\ket{64s_{1/2},m_{j}=1/2}$ of Cs and |$\ket{r}_{\mathrm{m}}\equiv\ket{60s_{1/2},m_{j}=1/2}$ of Rb that we now consider. Instead, we assume a two-photon optical Raman transition between the ground-state hyperfine sublevels |$\ket{0}_{\rm m}\equiv\ket{5s_{1/2},f=1,m_{f}=0}$ and |$\ket{1}_{\rm m}\equiv\ket{5s_{1/2},f=2,m_{f}=0}$ via the intermediate non-resonant state |$\ket{5p_{1/2},f=1,m_{f}=1}$. Note that for atoms in traps with the motional degrees of freedom cooled to $1-10\>\mu$K temperatures, the Doppler shift on the optical transition $\Delta_{\mathrm{Doppler}}\simeq 2\pi\times 10-40\leavevmode\nobreak$kHz is much smaller than the detuning $\Delta_{r}$ of the Rydberg dressing laser and can therefore be safely neglected. Simultaneously, the Doppler shift on the Raman transition between the ground-state sublevels is vanishingly small and does not affect the bowler-hat shaped transition spectrum of the Rb atoms when the Raman lasers are copropagating.
@@
where $\phi_{0}=2.03\>$rad is the single-atom dynamical phase accumulated during the transfer. We simulate the dynamics of the complete system, composed of one control qubit and $k=1,2,3,4$ target qubits, governed by the Hamiltonian (2). The resulting fidelity $\mathcal{F}=|\braket{\mathrm{GHZ}}{\Psi(\tau)}|^{2}$ of the GHZ (or Bell, for $k=1$) state is shown in the inset of Fig. 7. We obtain $\mathcal{F}\geq 0.98$ for all $k\leq 4$. The main factor that reduces the fidelity is the decay of the Rydberg state |$\ket{r}_{\mathrm{a}}$ of the Cs atom during the long gate time $\tau$. Then the fidelity decreases with increasing the number of target atoms due to the decay of the Rydberg state |$\ket{r}_{\mathrm{m}}$ with which we dress the qubit state |$\ket{1}_{\mathrm{m}}$ and therefore lose during the transfer some population from that state, $\Delta\braket{\sigma_{11}^{(j)}}\lesssim\frac{1}{2}\gamma_{r}\tau\frac{|\Omega_{r}|^{2}}{4\Delta_{r}^{2}}$. Again, going to higher Rydberg states for both Cs and Rb, and/or using the longer-lived $nP$ states, will increase the gate fidelity, and multiqubit Rydberg gates with error probability of $<1\%$ appear feasible under realistic experimental conditions. We note finally that in our simulations we neglected dephasing of the Rydberg state of Cs atoms and qubit states of the Cs and Rb atoms due to, e.g., external electric or magnetic field noise or laser phase fluctuations, since these are technical imperfections that should be suppressed in any experiment intended to achieve high fidelity quantum gates with atomic qubits.
@@ V Conclusions
Note added: After submission of this work related results appeared in [52].
@@ Appendix A Rydberg state parameters and calculation methodology
Atomic pair state interaction strengths were calculated with ARC [55, 54]. Pair states were calculated in a $\theta=\frac{\pi}{2},\phi=0$ geometry, with $m_{j}=1/2$. The interaction Hamiltonian was constructed from a range of states specified by $\Delta n=\pm 2,\Delta l=\pm 2$. For each Hamiltonian (Rb-Rb, Cs-Cs, and Rb-Cs), 250 eigenstates were extracted for the inter-atomic separation distances $r=1.5-6\;\mu$m. For each case, the arithmetic mean of the pair interaction curves, weighted by the overlap of the eigenstate with the target state $g=|\braket{\psi_{\rm pair-state}}{\psi_{\rm target}}|^{2}$, was fitted by a function of the form $V(r)=\frac{C_{3}}{r^{3}}+\frac{C_{6}}{r^{6}}$. Only states with $g\geq 0.4$ for Rb-Rb and Cs-Cs and $g\geq 0.01$ for Rb-Cs were included for fitting. The fits were performed over the interval $3\leavevmode\nobreak\ \mu$m $\leq r\leq 6\leavevmode\nobreak\ \mu$m, and results are presented in Fig. 8 and Table 2. For visual clarity, discrete points in Fig. 8 with $g<0.05$ are not shown.